Present: Rich Bennett, Bill Covey, Destry Fudge, Debra Harris, Winston Harris, Sharon Henry, Bobbie Parker, Suzy Shaw, Leota WhitfieldReview of Minutes of 12/1/97
- No changes.
- Martha has requested that the minutes (hereafter to be notes) of the Systems Staff meetings be published on the web.
- Room 10 move - Carpet was Murdochís product. Walls were to be up last week, need some magic gizmo to hook up door frames, didnít get it. Desperately trying to fix by the 15th - certified forward funds were used to purchase it. Only 2 file cabinets, maybe some part of the shipment was lost. Phones - cabling has been run. Should move by the last half of this month. May need electrical.
- Announced at PSSC that Dale will directly supervise Facilities.
- Yuletide Luncheon for Systems - Olive Garden Wednesday, 10 December at noon.
Ongoing Systems Updates Equipment Received/Installed/Repaired
- Iona wants newer Macs for PK Yonge. SEís and SE30ís will be surveyed.
- 20 Cows are in route.
- Last 20 Cows had no application software - anticipate application sharing will be going by the end of this week. Have enough Office 95 on the shelf to handle them. University contract is the least expensive way to get the software, but the 3rd party time delay prevents the media from being obtained in a timely fashion.
- Quiet at the last meeting.
- Core development - interoperability testing has not been done. IBMís 8265 will not be in until 15 December and the CISCO loaner 5500 is not in. Have 2 IBM 8274ís , not in the plan. With CISCO on the outer edge, Systems could use Fast Ethernet or ATM - will probably go with Fast Ether and then switch to ATM. Fast Ether is cheaper and has more bandwidth. Switches will be in after 1 January. Will then retire the routers - they could be used for white noise generator.
- Fiber cut at Dover seems to have been repaired. Adequate over the weekend. Home dial-up is faster than the Campus 10M.
- Microsoft does ping attacks and are the most popular web site.
- Application sharing.
- MSL - networking workstation - Bill will work on it.
- Documentation - being worked on.
- Bill found an oddity on 3rd floor Resource Services. Mac worked successfully, but a Cow from the same location wouldnít. Pole plug wasnít plugged into the hub.
- AFA baby web server - they came to get the new one. Trouble accessing the zip drive, wonít talk to the jazz drives. They must bring the old one back. AFA and Special Collections are the only ones who have given Rich the pointers.
- Memory was installed in the Directorsí suite.
- Library Office staff will be down to Chronda. Donna Marie is looking for a part-time position and Lana has left. Murdoch is gone from Facilities.
- Destry has a meeting with John to plan out the horizontal conduit run. Wiring closets in the north tower out. Barbara Oliver needs to do the budget. Router needs to be moved from the 2nd north tower to the closet. Multi-fiber pair is the hard part.
- Education has been underway since last week.
- LAC estimates are in. Sole source was killed by Purchasing. TDI subcontracted under Perry Parrish may squelch the have to bid.
- LAC is moving back to East over break. Phones are not in yet, just the estimates for them.
- $14,000 PO has been cut for Special Collections
- Charges are $7,000 for 5 phones - old type, not even modular jacks. Wiring is being put down which is sufficient to handle voice. Need clearance, PBX to handle this. Times have changed, possibly submit idea to upgrade the phones to the Davis Productivity Awards - new and efficient would be an enhancement and savings. Part of the problem may be the number of calls (trunk capacity).
- Destry installed an HP5P in the Reference area. It has been networked and added to NT.
- Sharon and Winston had a meeting with Laura from Personnel. She thinks her Cow isnít networked, but Destry had installed a baby hub and assigned the Cow an IP and informed Laura of it. She should work OK.
- Winston is not doing anything with Mellon.
- Accounting ILL migration - Winston is working on it.
- Library Information System - first project will be OPS tracking.
- Winston received a note from PhekSu - FCLA is coming to look at the spine labeling printing. Laurel will demonstrate on Thursday. We need to have access to the source code when it gets distributed unless FCLA will do 100% of the support and not refer to us.
- Destry out 12-17 December.
- Lee out Christmas week.
- Bill out Christmas week and New Year week (starting the 22nd and ending the 4th.
- Destry and Sharon out 26 December.
- Rich out 29 Dec - 2 Jan.
- Lee out 2 January.
Back to the Systems Homepage
Last updated March 7, 2000
by Debra Harris